IBM was mentioned in the chapter as having an uneven performance. Let’s check this out. Look at the stock chart and financial information below and answer the following questions. 1. How has IBM’s stock been doing recently? 2. Assuming IBM’s historical price-earnings ratio is 18, how does it currently stand? 3. Assuming its annual dividend yield is 2.5 percent, how does it currently stand? 4. Assuming IBM’s historical “LT” (long-term) debt/equity is 100 percent, how does it currently stand? Generally speaking, is it good or bad? 5. Assuming its historical return on assets is 10 percent, how does it currently stand? Generally speaking, is it good or bad? Please show work.

IBMs stock performance for today (Minimum 25 minutes delayed IBM Corporation (NYSE: IBM 11:16 AM EDT Jan 11, 2019 Last price Change Open Day high 52-week high 120.86 0.936076%) 121.58 121.62 171.13 Volume Previous close Day low 52-week low 973,053 121.79 120.20 105.94 The Stock chart page and related information is provided by West Corporation. Chart data is 25 minutes delayed. IBM Corporation does net maintain this page and is not responsible for the accuracy, completeness or timeliness of the information. The results are for illustrative purposes only and should not be relied on for investment purposes

Gross Margin (MRQ) C%) 46.93 Gross Margin (F Y) C%) 45.78 Legend Most Recent Quarter Mil Millions RTMA = Rolling Three Months Average TTM Trailing Twelve Months MRQ FY = Fiscal Year 1BM Corporation does not maintain this page and is not responsible for the accuracy, completeness or timeliness of the information. Historical financial information is provided by Thomson Reuters Corporation, a third party service. The company makes no representations or warranties with respect to the information contained herein and takes no responsibility for supplementing, updating, or correcting any such information.
Answer #1

1. How has IBM’s stock been doing recently?

IBM's stock has seen a decline recently (last three months) and hence it's performance can at the best be called "poor". The rationale for this conclusion is:

  • Please observe the price chart:
    • From a price level in excess of 160 in the month of January & February last year, it has come down to the vicinity of 120
    • A sharp decline can be seen in the month of October to November where price fell from 150 level to current level of 120
    • The stock never recovered fully since then
    • In the last three months, the prices had been in the close vicinity of 120 and hence volatility is quite low in this period.
  • Please look at the price and volume table:
    • 26 weeks (6 months) price change = - 16.73%
    • 52 weeks (1 year) price change = - 26.33%
  • 52 week high is 171.13 and 52 week low is 105.94
    • Current price of 121 is far more closer to 52 week low than to 52 week high

So, overall the price performance had been poor recently.

2. Assuming IBM’s historical price-earnings ratio is 18, how does it currently stand?

Current price, P = 120.69 (Recent price data under Price & Volume table)

EPS (TTM) = 12.25 (From Per share data table)

hence, P/E ratio = 120.69 / 12.25 = 9.85

EPS (MRQ) = 2.94 (From Per share data table)

Since this is EPS corresponding to a quarter, we need to annualise it. Annualised EPS = 4 x quarterly EPS = 4 x 2.94 = 11.76

P/E ratio = 120.69 / 11.76 = 10.26

Historical P/E = 18

  • P/E ratio based on TTM = 9.85; nearly half of historical average of 18
  • P/E ratio based on MRQ = 10.26; nearly 40% down to historical average

IBM's P/E has seen a significant decline over time.

3. Assuming its annual dividend yield is 2.5 percent, how does it currently stand?

Refer the table Dividend Yield:

  • Yield = 5.20%.
  • The same can also be calculated as per information below:
    • Annual dividend, D = 6.28;
    • Current market price, P = 120.69;
    • So dividend yield = D / P = 6.28 / 120.69 = 5.20%

The dividend yield has improved and nearly doubled. However the same can be partly be explained because of decline in price. The price has fallen over 26% in last one year and hence yield has improved. But this is part explanation. The other reason for the improvement yield may be the increase in absolute dividend per share.

4. Assuming IBM’s historical “LT” (long-term) debt/equity is 100 percent, how does it currently stand? Generally speaking, is it good or bad?

  • Observe the table "Financial Strength"
    • Long term debt to equity (MRQ) = 181.9%
    • Long term debt to equity (FY) = 226.42%
    • The ratio has worsened in comparison to its historical average.
    • Now, IBM has higher proportion of debt in its capital structure.

Generally speaking, the stock has worsened with respect to its historical performance. Further companies in the space of Information Technology, Consulting etc are nearly debt free and have near zero debt to equity ratio.So, in general, it's bad.

5. Assuming its historical return on assets is 10 percent, how does it currently stand?

  • Observe the table "Management Effectiveness"
    • Return on Assets (MRQ) = 8.86
    • Return on Assets (FY) = 9.25
    • Return on Assets (TTM) = 9.27
    • The ratio has worsened in comparison to its historical average. However the decline is limited to 10%.

So, generally speaking, this particular ratio has also declined but not sharply. But, overall it' still bad.